# GCC-Spam: Spam Detection via GAN, Contrastive Learning, and Character Similarity Networks

### v1

@@ 3.2.3 Similarity Loss Regularization
To ensure character substitutions remain plausible, the generator incorporates a character-level similarity loss using the LZASD similarity matrix shared with the discriminator. For each replaced position, a similarity score between the original and the selected substitute is computed. The similarity loss encourages the model to select visually and phonetically similar substitutes:
@@ 3.3 Discriminator
The baseline discriminator in our proposed framework adopts the architecture of the LZASD method, as introduced in [Original Paper Citation], specifically designed to tackle spam text detection by addressing adversarial deformations employed by spammers. This approach leverages character-level features to capture subtle patterns, such as deformed characters, that traditional word-based models often miss. The LZASD model processes a user message denoted as , where represents the -th character, aiming to predict the probability that the message is spam. Below, we detail the core components, their principles, and mathematical formulations, highlighting how LZASD effectively handles spam detection challenges.
The cornerstone of the LZASD framework is the character similarity network, constructed over a Chinese character repository , where denotes the -th character in the dictionary. This network quantifies similarities between characters based on two primary dimensions: glyph (e.g., stroke structure, such as left-right or up-down configurations) and pronunciation (e.g., pinyin components like initials, finals, and tones). Spammers often replace characters with visually or phonetically similar ones to evade detection, and this network aims to capture such adversarial substitutions. The overall similarity between two characters and is defined as
@@
To address the limitation of traditional word embeddings, which fail to recognize novel deformed words, LZASD generates character embeddings that incorporate semantic and emotional features. For a character , the embedding is computed as a weighted sum of embeddings of similar characters within the similarity network:
@@
The LZASD architecture excels in detecting adversarial spam by leveraging character similarity networks and self-attention mechanisms, making it a suitable baseline for our work. We extend this framework with [your novel contributions, e.g., contrastive learning, data augmentation] to address remaining challenges and enhance detection performance.
@@ 3.4 Contrastive Learning
To further enhance the discriminative power of the discriminator, we incorporate supervised contrastive learning (InfoNCE) into the embedding space. Let
@@ 4 Experiments
This section presents and analyzes the experimental results of the proposed GCC-Spam model. We compare our approach against three baselines: a LZASD architecture discriminator without generation or contrastive learning (Baseline), a spam detector enhanced by GAN-based adversarial training, and two data augmentation strategies (static and dynamic). All models are evaluated on a Chinese spam text dataset containing character-level and semantic-level perturbations. Metrics include precision, recall, F1-score, and overall accuracy.
@@ 4.1 Performance Comparison
| Class | Precision | Recall | F1-score | Support |
@@
| Class | Precision | Recall | F1-score | Support |
@@
| Class | Precision | Recall | F1-score | Support |
@@
| Class | Precision | Recall | F1-score | Support |
@@ 6 Enhanced Discriminator with Data Augmentation
The baseline discriminator, as described earlier, effectively captures adversarial deformations in spam text detection. However, its performance can be limited by the scarcity and lack of diversity in training data, particularly when facing novel or varied spam patterns. To address this, we propose an enhanced version of the discriminator by incorporating data augmentation, improving robustness and generalization. This appendix details the data augmentation approach, its principles, benefits, and integration into the training process.
@@
The training process integrates augmented data as follows. For each epoch, if an augmented data file exists, it is read and parsed into spam and normal samples. These augmented samples are accumulated across epochs to enrich the dataset. In each batch, a portion of samples is randomly drawn from augmented data, prioritizing these to introduce diversity, while the remainder comes from original training data, maintaining a balance between spam and normal texts. The batch is further refined to ensure balanced representation of classes using a pairing mechanism. Sentence embeddings are generated via self-attention, and the model computes a combined loss. The model is optimized using the Adam optimizer, and performance metrics (training loss, accuracy, test loss, and accuracy) are tracked. After each epoch, new augmented data is generated via the LLM, saved, and incorporated into subsequent training.
The benefits of this data augmentation approach are multifaceted. First, it significantly expands the training dataset, introducing diverse samples that simulate adversarial tactics, such as character substitutions, thereby improving the model’s ability to generalize to unseen spam patterns. Second, by focusing on misclassified samples, it targets the model’s weaknesses, enhancing robustness against challenging cases. Third, the iterative integration of augmented data across epochs allows continuous learning from varied examples, reducing overfitting to the original dataset. However, challenges include dependency on the quality of the LLM-generated samples, potential introduction of noise, and increased computational cost due to API calls. This enhanced discriminator, building on the LZASD foundation, leverages data augmentation to achieve superior spam detection performance, forming a critical component of our improved framework.

### v2

@@ 3.2.3 Similarity Loss Regularization
To ensure character substitutions remain plausible, the generator incorporates a character-level similarity loss using the character similarity matrix shared with the discriminator. For each replaced position, a similarity score between the original and the selected substitute is computed. The similarity loss encourages the model to select visually and phonetically similar substitutes:
@@ 3.3 Discriminator
The discriminator in our proposed framework adopts the architecture of the LZASD method, as introduced in [ref12], specifically designed to tackle spam text detection by addressing adversarial deformations employed by spammers. This approach leverages character-level features to capture subtle patterns, such as deformed characters, that traditional word-based models often miss. The discriminator processes a user message denoted as , where represents the -th character, aiming to predict the probability that the message is spam. Below, we detail the core components, their principles, and mathematical formulations, highlighting how the discriminator effectively handles spam detection challenges.
3.3.1 Character Similarity Network
The cornerstone of the discriminator is the character similarity network, constructed over a Chinese character repository , where denotes the -th character in the dictionary. This network quantifies similarities between characters based on two primary dimensions: glyph (e.g., stroke structure, such as left-right or up-down configurations) and pronunciation (e.g., pinyin components like initials, finals, and tones). Spammers often replace characters with visually or phonetically similar ones to evade detection, and this network aims to capture such adversarial substitutions. The overall similarity between two characters and is defined as
@@ 3.3.1 Character Similarity Network
3.3.2 Character Embedding
To address the limitation of traditional word embeddings, which fail to recognize novel deformed words, the discriminator generates character embeddings that incorporate semantic and emotional features. For a character , the embedding is computed as a weighted sum of embeddings of similar characters within the similarity network:
@@ 3.4 Contrastive Learning
To further enhance the discriminative power of the discriminator, we incorporate supervised contrastive learning (InfoNCE [ref23]) into the embedding space. Let
@@ 4 Experiments
This section presents and analyzes the experimental results of the proposed GCC-Spam model. We compare our approach against three baselines: a LZASD architecture discriminator without GAN or contrastive learning (Baseline), a spam detector enhanced by GAN-based adversarial training, and two data augmentation strategies (static and dynamic, which will be introduced in Appendix 6). All models are evaluated on a Chinese spam text dataset containing character-level and semantic-level perturbations. Metrics include precision, recall, F1-score, and overall accuracy.
@@ 4.1 Performance Comparison
| Class / Metric | Precision | Recall | F1-score | Support |
@@
| Class / Metric | Precision | Recall | F1-score | Support |
@@
| Class / Metric | Precision | Recall | F1-score | Support |
@@
| Class / Metric | Precision | Recall | F1-score | Support |
@@ 6 Enhanced Discriminator with Data Augmentation
The LZASD architecture discriminator, as described earlier, effectively captures adversarial deformations in spam text detection. However, its performance can be limited by the scarcity and lack of diversity in training data, particularly when facing novel or varied spam patterns. To address this, we propose an enhanced version of the discriminator by incorporating data augmentation, improving robustness and generalization. This appendix details the data augmentation approach, its principles, benefits, and integration into the training process.
@@
The training process integrates augmented data as follows. In the dynamic augmentation process, for each epoch, if an augmented data file exists, it is read and parsed into spam and normal samples. These augmented samples are accumulated across epochs to enrich the dataset. In each batch, a portion of samples is randomly drawn from augmented data, prioritizing these to introduce diversity, while the remainder comes from original training data, maintaining a balance between spam and normal texts. The batch is further refined to ensure balanced representation of classes using a pairing mechanism. Sentence embeddings are generated via self-attention, and the model computes a combined loss. After each epoch, new augmented data is generated via the LLM, saved, and incorporated into subsequent training.
In addition, static data augmentation is employed before training starts, with LLM API called to generate a large number of synthetic samples that mimic both spam and normal texts. These augmented samples are incorporated directly into the training dataset, with no distinction made between the original and augmented samples. This method effectively expands the training dataset without needing to differentiate between the two types of samples, allowing the model to train on a more diverse set of data from the beginning.
The benefits of data augmentation approach are multifaceted. First, it significantly expands the training dataset, introducing diverse samples that simulate adversarial tactics, such as character substitutions, thereby improving the model’s ability to generalize to unseen spam patterns. Second, by focusing on misclassified samples, it targets the model’s weaknesses, enhancing robustness against challenging cases. Third, the integration of augmented data allows continuous learning from varied examples, reducing overfitting to the original dataset. However, challenges include dependency on the quality of the LLM-generated samples, potential introduction of noise, and increased computational cost due to API calls. This enhanced discriminator leverages data augmentation to achieve superior spam detection performance, forming a comparison with our GCC-Spam method, and the experimental results are in the Section 4.